Force = 2.11 N i -3.19 N k acts on a pebble with position vector = 0.51 m j -2.12 m k relative to the origin.

What is the resulting torque on the pebble about the origin? Enter i, j, and k component.

What is the resulting torque on the pebble about the point (2.02 m, 0, -3.25 m)? Enter i, j, and k component.
Answer #1


Torque T = F x r

T = = F x r = i j k

2.11 0 -3.19

0 0.51 -2.12


so T = i ( 0 * -2.12 + 3.19 * 0.51 ) - j ( 2.12 * 2.11 -3.19 * 0) + k(2.11* 0.51 -0)

T 1.626 i - 4.47 j + 1.076 k

---------------------------------

change in posiiton
r = r2 -r1

r = 2.02 i - 3.25 k - 0.51j - 2.12 k

r = 2.02 i - 0.51 j + 5.37 k

so Torque T = F x r

T = i j k

2.11 0 -3.19

2.02 0 .51 5.37


T = i(0* 5.37 + 0.51* 3.19) - j (2.11 * 5.37 + 2.02 * 3.19) + k (2.11* 0.51 -0)

T = 1.626 i - 17.78 j + 1.07 k